Circular No. 6566

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) D. C. Lis, M. Gardner, and T. G. Phillips, Caltech Submillimeter Ob-

servatory (CSO); D. Bockelée-Morvan, N. Biver, J. Crovisier, and P. Colom, Observatoire de Paris; and D. Despois, Observatoire de Bordeaux, report: “Using the CSO on Feb. 16–20, we detected the 11(0,11)–10(0,10) 241.7-GHz, 12(0,12)–11(0,11) 263.7-GHz, and 16(0,16)–15(0,15) 351.6-GHz transitions of HNCO with integrated line areas on a T_{b} scale of 0.27 ± 0.04, 0.18 ± 0.02, and 0.46 ± 0.08 K km/s, respectively. The J(24–23) 218.3-GHz and J(29–28) 263.8-GHz transitions of HCCCN were detected with line areas of 0.24 ± 0.03 and 0.19 ± 0.03 K km/s. The estimated production rates for HNCO and HCCCN are 2.3 and 0.8 × 10^{27} molecules/s, respectively. The HCCCN/HCN abundance ratio is ∼ 0.06. The J(4–3) lines of HNC at 362.6 GHz and HCN at 354.5 GHz were detected with line areas of 5.8 ± 0.1 and 18.8 ± 0.1 K km/s, respectively, which corresponds to an HNC/HCN relative abundance of ∼ 0.25; on 1996 Dec. 5, we measured at the CSO a much lower ratio (0.05), similar to that observed in C/1996 B2 (Irvine et al. 1996, Nature 382, 418). We have also detected the J(3-2) lines of H^{13}CN at 259.0 GHz and HCN at 265.9 GHz, with line areas of 0.27 ± 0.02 and 23.0 ± 0.08 K km/s, respectively. This corresponds to an HCN/H^{13}CN abundance ratio of 90 ± 15. The first identifications of HNCO and H^{13}CN in a comet were obtained at the CSO of C/1996 B2 on 1996 Mar. 22–25, (Lis et al. 1997, Icarus, submitted).”

H. Matthews, Joint Astronomy Centre, Hilo, and Herzberg Institute of Astrophysics; and D. Jewitt, University of Hawaii, report submillimeter continuum measurements using the James Clerk Maxwell Telescope (+ B3 heterodyne SIS mixer receiver with a continuum backend). Continuum flux densities were measured at 344 GHz within a 14′′.5 beam; preliminary values on Feb. 9 and 16 UT were 329 ± 96 and 464 ± 62 mJy, respectively. The effective blackbody cross-section, ∼ 30 000 km^{2}, is 40 times larger than in comet 1P/Halley at similar heliocentric distances (e.g., Jewitt and Luu 1992, Icarus, 100, 187). Comparison with the results by Kreysa et al. (IAUC 6555) shows that the dust cross-section is increasing faster than would be expected on the basis of the decreasing heliocentric distance alone.

Naked-eye m_{1} estimates: Feb. 18.18 UT, 1.1 (K. Sárneczky, Szeged, Hungary); 20.20, 1.0 (R. J. Bouma, Groningen, The Netherlands); 22.49, 0.9 (W. Dillon, Missouri City, TX); 23.47, 0.8 (G. W. Kronk, Troy, IL); 24.42, 0.8 (D. W. E. Green, Lexington, MA).

1997 February 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green